OpenStudy (anonymous):

3^(x – 4) = 7^(x + 9) please

OpenStudy (turingtest):

I'd like to gently remind all those who would attempt to answer this question that it against the code of conduct to post the answer without sufficient explanation. Thank you.

OpenStudy (anonymous):

thanks alan

OpenStudy (anonymous):

alan turing

OpenStudy (anonymous):

log both sides

OpenStudy (anonymous):

log3^(x-4)=log7^(x+9)

OpenStudy (anonymous):

remember the exponent rule?

OpenStudy (anonymous):

(x-4)log3 = log3^(x-4)

OpenStudy (anonymous):

so (x-4)log3=(x+9)log7

OpenStudy (anonymous):

...thinking

OpenStudy (anonymous):

expand, collect like terms

OpenStudy (anonymous):

o got doff answers but none of them are x ≈ –25.86 x ≈ –2.59 x ≈ –0.39 x ≈ –0.04

OpenStudy (anonymous):

then either you're wrong or the answers are wrong

OpenStudy (turingtest):

what answer did you get? could you show us how you got it please?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i am using diff calculater sights and iv got answers 0 and 1

OpenStudy (turingtest):

what did you put into your calculator?

OpenStudy (anonymous):

-2.59? looks like -25.85589026...

OpenStudy (anonymous):

maybe it's a misprint of the decimal place? lol

OpenStudy (anonymous):

ok thank you thats what i wanted to know

OpenStudy (anonymous):

yah maybe

OpenStudy (anonymous):

what are the probabilities of you getting it wrong given that i've gotten it wrong and the question misprinting the answer?

OpenStudy (anonymous):

and another one, the probability of me getting it wrong given that you've gotten it wrong

OpenStudy (anonymous):